# Static Word Embeddings for Sentence Semantic Representation

### v1

@@ 4.2 Baselines
In monolingual experiments, we include two traditional SWEs as our baselines, namely fastText Bojanowski et al. (2017) and Sent2vec. Both models are self-supervised based on word co-occurrence information, and Sent2vec performs better on sentence-level tasks. We also compare our model against two recent GitHub-hosted models with no paper: WordLlama and Model2Vec. We use wordllama-l3-supercat for WordLlama, which is based on LLaMa 3, and potion-base-32M for Model2Vec, the latest and state-of-the-art model (released in January 2025) based on BGE-base. SWE models, including ours produce sentence vectors by averaging SWEs, followed by L2-normalisation.
In cross-lingual experiments, we compare our model against the multilingual Model2Vec model (M2V_multilingual_output). For English–German, we also evaluate cross-lingual SWEs available at the MUSE GitHub repository [^6] https://github.com/facebookresearch/MUSE Conneau et al. (2018), which align English and German fastText embeddings using a bilingual dictionary and have been widely used in the research community.
Models ($d$) $100\rho$ runtime (s) GPU CPU MiniLM-L6 (384) 85.4 1.7 8.8 GTE-base (768) 87.2 3.4 52.7 LLM2Vec (4,096) 88.1 30.3 $>$10,000 WordLlama (512) 81.5 0.4 Model2Vec (512) 80.8 0.3 OURS (256) 83.1 0.4
@@ 4.3 Experiments on English Tasks
We first evaluate monolingual (English) models on semantic textual similarity (STS), a well-established task for sentence semantic evaluation. It is a task of calculating textual similarity for each sentence pair, and models are evaluated based on Spearman’s rank correlation $\rho$ between their predictions and human judgements. Table 1 shows the results on STS12–17, 22, STS Benchmark (STS-B), SICK-R, and BIOSSES (BIO) data sets Agirre et al. (2012, 2013); Bandhakavi et al. (2014); Biçici (2015); Agirre et al. (2016); Chen et al. (2022); Marelli et al. (2014); Soğancıoğlu et al. (2017). The first five rows show performance of STs for reference (not as baselines of our model): SimCSE [^7] We use the supervised one, which performs better than the unsupervised one. is a pioneering ST, which often serves as a strong baseline in STS research; MiniLM-L6 is a popular light-weight ST based on MiniLM Wang et al. (2020); BGE and GTE are the models used to train Model2Vec and OURS; and LLM2Vec is a recent large model built on LLaMa-3-8B. The table shows that OURS substantially outperforms all SME baselines on most data sets, even with the smallest dimensionality ($d=256$). Our model also outperforms MiniLM on STS14 and STS16, and SimCSE on BIOSSES, where the sentences are sampled from biomedical domains, indicating the robustness of OURS to various types of text. Table 2 compares the scores and runtime on STS15. While falling behind STs in accuracy, OURS (and other SWEs) can run extremely fast on CPUs (20 times faster than MiniLM) without sacrificing the accuracy very much.
@@
Additionally, to further evaluate the effectiveness of our model on other tasks such as sentence classification, we also run experiments on Massive Text Embedding Benchmark (MTEB) Muennighoff et al. (2023). It consists of 56 data sets covering 7 different tasks, namely classification, pair classification, clustering, reranking, retrieval, STS, and summarisation evaluation. On each task, a model encodes text and returns its fixed-length vector, which is then used on downstream tasks (e.g. for calculating textual features or similarities). Models are evaluated using task-specific metrics (e.g. precision) on each data set. Since MTEB contains both sentence-level and document-level tasks, and our focus is on sentence semantic representation, we report two scores: Avg, the average across all 56 data sets; and Avg-s2s, the average across the 33 data sets labelled as “s2s”, where the input text is a sentence rather than a document. Table 3 presents the results, showing that our model substantially outperforms the other SWEs in Avg and Avg-s2s, even with the smallest embedding size. Moreover, our model outperforms SimCSE in both Avg and Avg-s2s, which is surprising given that our model does not use any sequential function like Transformer. Interestingly, we also find that ensembling OURS with WordLlama and Model2Vec further boosts performance (64.55 and 53.28 in Avg-s2s and Avg), which we describe in Appendix B. We also show the scores on each task in Table 10 in Appendix.
@@ 4.4 Experiments on Cross-Lingual Tasks
We first evaluate cross-lingual SWEs on translation retrieval, a task of retrieving the translation of the input text based on cosine similarity. Table 4 shows the results (F1 scores) on BUCC Zweigenbaum et al. (2017) and Tatoeba Artetxe and Schwenk (2019); community (2021) data sets for three language pairs: English–{German, Chinese, Japanese}. The first row shows the performance of mGTE, and the rest are cross-lingual SWEs. It clearly shows that OURS outperforms the SWE baselines by a large margin, and performs fairly well on all language pairs (e.g. 96.3% for English-German on BUCC). This suggests the applicability of our model as a light-weight method for extracting translation candidates, which could be further validated by computationally expensive models such as mGTE. Lastly, we also evaluate SWEs on cross-lingual STS (STS17 and STS22), a task of measuring textual similarity across two different languages, and Table 5 presents the results. Again, OURS is the best performing SWE, demonstrating its effectiveness on cross-lingual tasks.
@@ 5 Analysis
5.1 Ablation Studies
We evaluate our SWEs ($d=256$) obtained in each section and Table 6 shows the results. The column under “en” shows performance on MTEB (Avg), indicating that sentence-level PCA and knowledge distillation substantially improve the SWEs extracted in Section 3.1. Other important findings that are omitted from Table 6 include: (1) applying word-level PCA in Section 3.2 performs worse than sentence-level PCA (48.1 vs. 49.9); and (2) performing knowledge distillation without PCA brings no improvement (44.4) [^8] Performance increases on some tasks but drops on others., indicating that parameter initialisation is a key to success in knowledge distillation. The last three columns in Table 6 show the average scores on BUCC and Tatoeba for each language pair, indicating that PCA and contrastive learning in Section 3.3 are very effective on cross-lingual tasks. It also reveals that ABTT, i.e. removing the first $r$ PCs from SWEs, plays a vital role in generating cross-lingual SWEs.
Table 7 shows performance of our models trained with different STs, namely SimCSE, Nomic Nussbaum et al. (2024), and GTE-{small/base/large}. It indicates that OURS performs better using better STs (SimCSE < Nomic < GTE) as its teacher model, suggesting that performance of our model might keep increasing as ST models improve. That being said, our model does not benefit from using the largest GTE model, possibly due to the large discrepancy of the parameter sizes between the teacher and student models during knowledge distillation. Performance on each MTEB task is shown in Table 12 in Appendix.
5.2 Effects of PCA
@@
To investigate why PCA with ABTT is particularly effective on cross-lingual tasks, we analyse what information is encoded in the first two PCs and find that it captures the variance largely based on the language identity. To visualise this, we randomly sample 1,000 language-specific words/tokens from English and Japanese and plot their embeddings based on the values in the first two PCs (PC1 and PC2). The result is presented in Figure 1 (the leftmost plot), which clearly shows that the embeddings are separated based on language. Given this finding, we also visualise the embeddings using t-SNE van der Maaten and Hinton (2008) before and after applying PCA wo/w ABTT, and the results are also shown in Figure 1. We can see that the embeddings create language clusters before PCA and after PCA without ABTT, but applying PCA with ABTT renders the embeddings more language-agnostic by removing the language-identity components. We also observe the same trend in English–Chinese/German, as shown in Figures 5 and 6 in Appendix. Such language-specific dimensions can act as noise when measuring semantic similarity across different languages, [^9] For instance, Japanese sentences with many English loanwords can have high similarities to English sentences regardless of the semantics. explaining why ABTT is crucial in cross-lingual SWEs. Furthermore, this observation makes us wonder whether such language segregation also occurs in the mGTE embedding space; we investigate this and find that it does create language clusters at lower layers of Transformer, but generates more language-agnostic embeddings at higher layers. Since it deviates from the main subject of this study, we discuss more details in Appendix A.
Lastly, we also analyse the dominant PCs of monolingual (English) SWEs, and find that some of them seem to capture linguistic styles or domains. Table 8 shows examples of the words that have the five largest and smallest values in PC1, PC2, and PC6 among the 10k most frequent words in $V$. In PC1 and PC2, words seem to be clustered based on the word frequency in casual text, whereas in PC6, date-related words seem to have large values. To verify whether PC1 and PC2 are actually related to textual formality, we conduct a simple experiment using the data set released by Chhaya et al. (2018), where 10 human annotators judge the formality of 960 emails. We encode each email by averaging SWEs and measure the Pearson correlation coefficient between the values in PC1, PC2, or PC6, and the mean scores of the 10 annotators. We observe that the correlations are: $-0.63$, $0.37$, and $-0.11$ for PC1, PC2, and PC6, confirming a moderate or weak correlation for the first two PCs. [^10] We surmise that this may be attributed to the use of sentences sampled from a web corpus in our model, which is a mixed bag of formal and casual text. It would be interesting to see how the selection of text data affects the outcome. Albeit those components capture the variance of sentences effectively, they would not necessarily serve as useful features on semantic tasks, explaining why PCA w/ ABTT works well for monolingual SWEs as well, [^11] Yet, ABTT often harms performance on classification. which Mu and Viswanath (2018) describe as counter-intuitive.
@@ 5.3 Analysis of Word Embedding Norm
Figure 2: Comparison of word embedding norms obtained in Sections 3.1, 3.2, and 3.3. For normalisation, norms of each POS tag are divided by the maximum value of all tags (i.e. the norm of Proper Noun).
To investigate how our method optimises word embeddings, we analyse the norms of SWEs obtained in Sections 3.1, 3.2, and 3.3; we denote them as $\|\mathrm{E}_{i}(w)\|$ ($i=1,2,3$). For the purpose of analysis, we compare the norms of the 10k most frequent words in Brown Corpus, which consists of text in various domains and provides part-of-speech (POS) tags for each word. First, we compare the mean norms of word embeddings for each POS tag, [^12] We use the most frequent tag for each word. and Figure 2 shows the results. For normalisation, we divide the norms of each POS tag by that of “Proper Noun”, which always has the largest value regardless of $i$. It shows that the differences in the norms are emphasised after PCA and knowledge distillation. In particular, the relative norms of “Article” (e.g. the), “Preposition” (e.g. of), and “Numeral” (e.g. 2025, first) drop sharply because these types of words do not affect sentence semantics very much. On the other hand, “Proper Noun” and “Noun” have the largest norms after PCA and knowledge distillation, as they often convey important information. This result is somewhat akin to the previous finding that ST tends to emphasise nominal information of a sentence Nikolaev and Padó (2023). Lastly, we find that our method induces a similar effect to SIF, a heuristic used in Model2Vec — we measure Spearman’s rank correlation between the norm and word frequency rank and find that it increases after PCA and knowledge distillation ($0.39\rightarrow 0.43\rightarrow 0.49$), meaning both steps decrease the norms of high-frequency words. Compared to SIF, however, knowledge distillation adjusts norms in a more nuanced way, considering not only word frequency (which is implicitly encoded by ST Kurita et al. (2023)) but also POS tags and influence of each word on sentence semantics.
@@ 6 Conclusion
We propose new word embeddings optimised for sentence semantic representation. We extract embeddings from Sentence Transformers and improve them with sentence-level PCA and knowledge distillation or contrastive learning. Our rigorous experiments show that our model outperforms baselines on both monolingual and cross-lingual tasks. Our analyses also reveal that our method removes embedding components that are not relevant to sentence semantics, and adjusts the vector norms based on the influence on sentence semantics.
@@ 7 Limitations
Since our model (and other SWEs) does not consider word order information (i.e. a bag-of-words model), its performance is limited on tasks where models are required to understand long context. Importantly, however, there is always a trade-off between model performance and computational complexity; in this work, we focus on improving the cost-efficiency, an important factor for real-world applications that is nonetheless largely neglected among the research community these days. To improve performance (at the cost of computational cost), one can simply add a small network that can capture sequential information (e.g. Transformer, CNN, LSTM Hochreiter and Schmidhuber (1997)) on top of our word embeddings, and train them with labelled data following the training process of Sentence-Transformer models. Another idea to better capture sentence semantics is to represent text with a set of word embeddings with positional information (rather than taking their average), and calculate textual similarity by measuring the distance (or optimal transport cost) between two distributions, e.g. Word Mover’s Distance Kusner et al. (2015).
Appendix A Analysis of mGTE Embeddings

### v2

@@ 4.2 Baselines
In monolingual experiments, we include two traditional SWEs as our baselines, namely fastText Bojanowski et al. (2017) and Sent2vec. Both models are self-supervised based on word co-occurrence information, with Sent2vec performing better on sentence-level tasks. We also compare our model against two recent GitHub-hosted models: WordLlama and Model2Vec. We use wordllama-l3-supercat for WordLlama, which is based on LLaMa 3, and potion-base-32M for Model2Vec, the latest and state-of-the-art model (released in January 2025) based on BGE-base. SWE models, including ours, produce sentence vectors by averaging SWEs and applying L2-normalisation.
In cross-lingual experiments, we compare our model against the multilingual Model2Vec model (M2V_multilingual_output). For English–German, we also evaluate the cross-lingual SWE available at the MUSE GitHub repository [^10] https://github.com/facebookresearch/MUSE Conneau et al. (2018), which aligns English and German fastText embeddings using a bilingual dictionary. This model has been widely used in research papers.
Models ($d$) $100\rho$ runtime (s) GPU CPU MiniLM-L6 (384) 85.4 1.7 8.8 GTE-base (768) 87.2 3.4 52.7 LLM2Vec (4,096) 88.1 30.3 >>10,000 WordLlama (512) 81.5 0.4 Model2Vec (512) 80.8 0.3 OURS (256) 83.1 0.4
@@ 4.3 Experiments on English Tasks
We first evaluate monolingual (English) models on semantic textual similarity (STS), a well-established NLP task for sentence semantic evaluation. It is a task of calculating textual similarity for each sentence pair, and models are evaluated based on Spearman’s rank correlation $\rho$ between their predictions and human judgements. Table 1 shows the results on STS12–17, 22, STS Benchmark (STS-B), SICK-R, and BIOSSES (BIO) data sets Agirre et al. (2012, 2013); Bandhakavi et al. (2014); Biçici (2015); Agirre et al. (2016); Chen et al. (2022); Marelli et al. (2014); Soğancıoğlu et al. (2017). The first five rows show performance of STs for reference (not as baselines of our model): SimCSE [^11] We use the supervised one, which performs better than the unsupervised one. is a pioneering ST, which often serves as a strong baseline in STS research; MiniLM-L6121212sentence-transformers/all-MiniLM-L6-v2 is a popular light-weight ST based on MiniLM Wang et al. (2020); BGE and GTE are the models used to train Model2Vec and OURS, respectively; and LLM2Vec is a recent large-scale model built on LLaMa-3-8B. The table shows that OURS substantially outperforms all SME baselines on most data sets, even though it has the smallest dimensionality ($d=256$). Our model also outperforms MiniLM-L6 on STS14 and STS16 and SimCSE on BIOSSES, where the sentences are sampled from biomedical domains, indicating the robustness of OURS to various types of text. Table 2 compares the scores and runtime on STS15. While falling behind STs in accuracy, OURS (and other SWEs) can run extremely fast (20 times faster than MiniLM on CPUs) without sacrificing the accuracy very much.
Additionally, to further evaluate the effectiveness of our model on other tasks such as sentence classification, we also run experiments on Massive Text Embedding Benchmark (MTEB) Muennighoff et al. (2023). It consists of 56 data sets covering 7 different tasks, namely classification, pair classification, clustering, reranking, retrieval, STS, and summarisation evaluation. On each task, a model encodes text and returns its fixed-length vector, which is then used on downstream tasks (e.g. for calculating textual features or similarities). Models are evaluated using task-specific metrics (e.g. precision) on each data set. Since MTEB contains both sentence-level and document-level tasks, and our focus is on sentence semantic representation, we report two scores: Avg, the average across all 56 data sets; and Avg-s2s, the average across the 33 data sets labelled as “s2s”, where the input text is a sentence rather than a document. Table 3 presents the results, showing that our model substantially outperforms the other SWEs both in Avg and Avg-s2s, even with the smallest embedding size. Moreover, our model outperforms SimCSE, which is surprising given that our model does not use any sequential function like Transformer. Interestingly, we also find that ensembling OURS with WordLlama and Model2Vec further boosts performance (64.55 and 53.28 in Avg-s2s and Avg), which we describe in detail in Appendix D. We also show the scores on each task in Table 13 in Appendix.
@@ 4.4 Experiments on Cross-Lingual Tasks
We first evaluate cross-lingual SWEs on translation retrieval, a task of retrieving the translation of the input text based on cosine similarity. Table 4 shows the results (F1 scores) on BUCC Zweigenbaum et al. (2017) and Tatoeba Artetxe and Schwenk (2019); community (2021) data sets for three language pairs: English–{German, Chinese, Japanese}. The first row shows the performance of mGTE, and the rest are cross-lingual SWEs. The table clearly shows that OURS outperforms the SWE baselines by a large margin, and performs fairly well on all language pairs (e.g. 96.3% for English-German on BUCC). This suggests the applicability of our model as a light-weight approach to extracting translation candidates, which could be further validated by computationally expensive models such as mGTE. Lastly, we also evaluate SWEs on cross-lingual STS (STS17 and STS22), a task of measuring textual similarity across two different languages, and Table 5 presents the results. Again, OURS is the best performing SWE, demonstrating its effectiveness on cross-lingual tasks.
It should be noted that mGTE and Model2vec are multilingual models (i.e. aligned across more than two languages) whereas MUSE and OURS are bilingual ones, and this can place our model in a favourable position in bilingual evaluation. Given this concern, we also try generating multilingual SWEs aligned across the four languages (English, German, Chinese, and Japanese) using our model, and find that it still outperforms the SWE baselines by a large margin; see Appendix A for details.
5 Analysis
5.1 Ablation Studies
@@
To verify the effectiveness of our method, we evaluate our SWEs ($d=256$) obtained in Sections 3.1, 3.2, and 3.3, and see whether the performance improves after each step. Table 6 shows the results. The column under “en” shows performance on MTEB (Avg), demonstrating that sentence-level PCA and knowledge distillation substantially improve the SWEs extracted in Section 3.1. Importantly, our model surpasses the performance of fastText (43.1) even without fine-tuning in Section 3.3, demonstrating the effectiveness of extracting SWEs from Sentence Transformers. Other important findings that are not presented in Table 6 (due to space restrictions) include: (1) applying word-level PCA in Section 3.2 performs worse than sentence-level PCA (48.1 vs. 49.9); and (2) performing knowledge distillation without PCA (i.e. ablating Section 3.2) results in poor performance (44.4) [^13] Performance increases on some tasks but drops on others., indicating that parameter initialisation is a key to success in knowledge distillation. The last three columns in Table 6 show the average scores on BUCC and Tatoeba for each language pair. The results indicate that the SWEs obtained in Section 3.2 already achieve good scores, and fine-tuning them in Section 3.3 further boosts performance. Another important observation is that applying ABTT in Section 3.2, i.e. removing the first $r$ PCs from SWEs, plays a vital role in generating cross-lingual SWEs; we will provide an in-depth analysis on its effect in Section 5.2.
Table 7 shows the performance of our models trained with different STs, namely SimCSE, Nomic Nussbaum et al. (2024), [^14] Nomic is trained to encode text with a task-specific prefix prompt, and hence we prepend the prompt ”classification: ” to every input text in Section 3.1 and Section 3.3. and GTE-{small/base/large}. It indicates that OURS performs better using better STs (SimCSE < Nomic < GTE) as its teacher model, [^15] Note that SimCSE uses [[CLS]] pooling while Nomic and GTE use mean pooling for generating sentence embeddings. While our model works reasonably well with both model types, this could be another reason why it performs better with Nomic and GTE than SimCSE, as our method extracts SWEs in Section 3.1 by averaging embeddings across $N$ sentences. suggesting that its performance might further increase as ST models evolve. That being said, our model does not benefit from using the largest GTE model, possibly due to the large discrepancy of the parameter sizes between the teacher and student models during knowledge distillation. Performance on each MTEB task is shown in Table 15 in Appendix.
@@ 5.2 Effects of PCA
To investigate why PCA with ABTT is particularly effective on cross-lingual tasks, we analyse what information is encoded in the first two PCs, and find that it captures the variance largely based on the language identity. To visualise this, we randomly sample 1,000 language-specific words/tokens from English and Japanese and plot their embeddings based on the values in the first two PCs (PC1 and PC2). The result is presented in Figure 1 (the leftmost plot), which clearly shows that the embeddings are separated based on language. Given this finding, we also visualise the embeddings using t-SNE van der Maaten and Hinton (2008) before and after applying PCA wo/w ABTT, and the results are also shown in Figure 1. We can see that the embeddings create language clusters before PCA and after PCA without ABTT, but applying PCA with ABTT renders the embeddings more language-agnostic by removing the language-identity components. [^16] We also confirm this finding by measuring the silhouette scores, which are: 0.10 before PCA; 0.11 after PCA w/o ABTT; and 0.02 after PCA w/ ABTT, respectively. We also observe the same trend in English–Chinese/German, as shown in Figures 5 and 6 in Appendix. Such language-specific dimensions can act as noise when measuring semantic similarity across different languages, [^17] For instance, Japanese sentences with many English loanwords can have high similarities to English sentences regardless of the semantics. explaining why ABTT is crucial in cross-lingual SWEs. Furthermore, this observation makes us wonder whether such language segregation also occurs in the mGTE embedding space; we investigate this and find that it does create language clusters at lower layers of Transformer, but generates more language-agnostic embeddings at higher layers. Since it deviates from the main subject of this study, we discuss more details in Appendix C.
@@
Lastly, we also analyse the dominant PCs of monolingual (English) SWEs, and find that some of them seem to capture linguistic styles or domains. Table 8 shows examples of the words that have the five largest and smallest values in PC1, PC2, and PC6 among the 10k most frequent words in $V$. In PC1 and PC2, words seem to be clustered based on the word frequency in casual text, whereas in PC6, date-related words seem to have large values. To verify whether PC1 and PC2 are actually related to textual formality, we conduct a simple experiment using the data set released by Chhaya et al. (2018), where 10 human annotators judge the formality level of 960 emails. We encode each email by averaging SWEs and measure the Pearson correlation coefficient between the values in PC1, PC2, or PC6, and the mean scores of the 10 annotators. We observe that the correlations are: 0.63-0.63, $0.37$, and 0.11-0.11 for PC1, PC2, and PC6, confirming a moderate or weak correlation for the first two PCs. [^18] We surmise that this may be attributed to the use of sentences sampled from a web corpus in our model, which is a mixed bag of formal and casual text. It would be interesting to see how the selection of text data affects the outcome. Albeit those components capture the variance of sentences effectively, they would not necessarily serve as useful features on semantic tasks, explaining why PCA w/ ABTT works well for monolingual SWEs as well. [^19] But on classification tasks, ABTT often harms performance.
Figure 2: Comparison of word embedding norms obtained in Sections 3.1, 3.2, and $3.$ For normalisation, norms of each POS tag are divided by the maximum value of all tags (i.e. the norm of Proper Noun).
@@ 5.3 Analysis of Word Embedding Norm
To investigate how our method optimises word embeddings, we analyse the norms of SWEs obtained in Sections 3.1, 3.2, and 3.3; we denote them as $\|\mathrm{E}_{i}(w)\|$ ($i=1,2,3$). For the purpose of analysis, we compare the norms of the 10k most frequent words in Brown Corpus, which consists of text from various domains and provides part-of-speech (POS) tags for each word. First, we compare the mean norms of word embeddings for each POS tag; for words that are annotated with multiple POS tags (e.g. notice), we use the most frequent one. The results are shown in Figure 2. For normalisation, we divide the norms of each POS tag by that of “Proper Noun”, which always has the largest value regardless of $i$. It shows that the differences in the norms are emphasised after PCA and knowledge distillation. In particular, the relative norms of “Article” (e.g. the), “Preposition” (e.g. of), and “Numeral” (e.g. 2025, first) drop sharply, likely because these types of words do not affect sentence semantics very much. On the other hand, “Proper Noun” and “Noun” have the largest norms after PCA and knowledge distillation, as they often convey information relevant to semantics. This result is somewhat akin to the previous finding that ST tends to emphasise nominal information of a sentence Nikolaev and Padó (2023). In Appendix B, we also perform the same analysis for SWE baselines (fastText, WordLlama, and Model2Vec) and reveal that fastText assigns relatively large norms to articles, suggesting it is not well optimised for sentence representation.
Lastly, we find that our method induces a similar effect to SIF, a heuristic used in Model2Vec — we measure Spearman’s rank correlation between the norm and word frequency rank and find that it increases after PCA and knowledge distillation ($0.39\rightarrow 0.43\rightarrow 0.49$), meaning both steps decrease the norms of high-frequency words. Compared to SIF, however, knowledge distillation adjusts norms in a more nuanced way, considering not only word frequency (which is implicitly encoded by ST Kurita et al. (2023)) but also POS tags and influence of each word on sentence semantics.
@@ 6 Conclusion
We propose new word embeddings optimised for sentence semantic representation. We extract embeddings from Sentence Transformers and improve them with sentence-level PCA followed by knowledge distillation or contrastive learning. Our rigorous experiments show that our model outperforms baselines on both monolingual and cross-lingual semantic tasks. Our in-depth analyses on word embeddings also reveal that our method successfully removes embedding components that are not very relevant to sentence semantics, and adjusts the vector norms based on the influence on sentence semantics.
@@ 7 Limitations
Since our model (and other SWEs) does not consider word order information (i.e. a bag-of-words model), its performance is limited on tasks where models are required to understand long context, such as document retrieval. Importantly, however, there is always a trade-off between model performance and computational complexity; in this work, we focus on improving the cost-efficiency, an important factor for real-world applications that is nonetheless largely neglected among the research community these days. To improve performance (at the cost of computational cost), one can simply add a small network that can capture sequential information (e.g. Transformer, CNN, LSTM Hochreiter and Schmidhuber (1997)) on top of our word embeddings, and train them with labelled data as performed during the training process of Sentence-Transformer models. Another idea to better capture sentence semantics is to represent text with a set of word embeddings with positional information (rather than taking their average), and calculate textual similarity by measuring the distance (or optimal transport cost) between two distributions, e.g. Word Mover’s Distance Kusner et al. (2015).
Appendix A Results on Multilingual Embeddings
In addition to bilingual SWEs, we also try generating multilingual SWEs using our model, where English, German, Chinese, and Japanese are all embedded in the same space (as in Model2Vec). To achieve this, we apply the sentence-level PCA to the concatenation of the sentence vectors sampled from each language (i.e. setting $L$ to $4$ in Section 3.2). For multilingual SWEs, we skip the fine-tuning step in Section 3.3 because it is specifically designed for bilingual training; however, note that it could be easily extended to multilingual training by jointly optimising the bilingual loss $\mathcal{L}_{CL}$ for multiple language pairs. Table 10 and Table 11 compare the performance of our bilingual and multilingual SWEs. Although the multilingual model performs worse than the bilingual one (which benefits from fine-tuning), it still substantially exceeds the performance of the SWE baselines (MUSE and Model2Vec) shown in Table 4 and Table 5.
Appendix B Analysis of Baseline Embedding Norms
POS OURS FT WL M2V Article 0.37 0.55 0.22 0.24 Preposition 0.53 0.66 0.33 0.57 Adverb 0.71 0.72 0.52 0.85 Numeral 0.85 0.76 0.66 0.79 Verb 0.79 0.81 0.67 0.86 Adjective 0.87 0.81 0.74 0.90 Noun 0.91 0.86 0.84 0.91 Proper Noun 1.00 1.00 1.00 1.00
Table 12: Comparison of word embedding norms generated by fastText (FT), WordLlama (WL), Model2Vec (M2V), and our model (OURS). For normalisation, norms of each POS tag are divided by the maximum value of all tags, which is the norm of Proper Noun for all models.
Table 12 compares word embedding norms generated by fastText (FT), WordLlama (WL), Model2Vec (M2V), and our model (OURS). Each value is normalised (divided) by the norm of Proper Noun within each model, as done in Figure 2. Interestingly, it shows that Proper Noun has the largest norm for all models. Looking at each model, fastText has relatively large norms for Article and Preposition (which usually have a small influence on sentence semantics), suggesting it is not well optimised for sentence representation (as we demonstrated in our experiments). For WordLlama, only Noun and Proper Noun have larger values than 0.8, suggesting it emphasises nominal information more than the other models. This partially explains why it performs well on document retrieval tasks on MTEB (as shown in Table 13), where capturing topical information is arguably more important than the semantics. Model2Vec and OURS display similar patterns, with Model2Vec assigning a higher weight to Verb and Adverb and OURS emphasising Numeral more.
Appendix C Analysis of mGTE Embeddings
Figure 3: The t-SNE visualisation of sentence embeddings produced by mGTE-base at the 1st, 4th, 8th, and 12th (= last) layers of Transformer for English (blue) and Japanese (orange) sentences. The encoded sentences are 1,000 pairs of translations sampled from a parallel corpus.
Figure 4: The t-SNE visualisation of sentence embeddings produced by mGTE-MLM-base (the backbone masked language model of mGTE before fine-tuning) at the 1st, 4th, 8th, and 12th (= last) layers of Transformer for English (blue) and Japanese (orange) sentences. The encoded sentences are 1,000 pairs of translations sampled from a parallel corpus.
Figure 5: Scatter plots of our cross-lingual SWEs in English (blue) and Chinese (orange). The leftmost shows the 1st and 2nd principal components, and the rest are t-SNE visualisation before and after applying PCA with/without ABTT.
Figure 6: Scatter plots of our cross-lingual SWEs in English (blue) and German (orange). The leftmost shows the 1st and 2nd principal components, and the rest are t-SNE visualisation before and after applying PCA with/without ABTT.